OpenStudy (anonymous):

can somebody please show me how to solve this? convert the angle measure to radians to the nearest hundredth of a radian 32°

OpenStudy (anonymous):

Oh, I've never heard of anything like this. What math class are you taking?

OpenStudy (anonymous):

trig :/

OpenStudy (anonymous):

it really confuses me and its really hard to learn online

OpenStudy (anonymous):

i dont even have a textbook haha

OpenStudy (phi):

there are pi radians in 180 degrees multiply degrees by \[ \frac{ \pi \text{ radians}}{180º} \]

OpenStudy (phi):

make sure to use enough digits in pi (use a calculator, or the number 3.14159)

OpenStudy (anonymous):

how could i put that in a calculator?

OpenStudy (phi):

calculators have a key for pi. Look for it.

OpenStudy (anonymous):

yes but it doesnt have a radians button.. haha

OpenStudy (whpalmer4):

\(2\pi\) radians = \(360\) degrees or written as a fraction, \[\frac{radians}{degrees} = \frac{2\pi}{360}\] Cross-multiply to find the one you don't know. I prefer remembering the whole circle to half the circle, which is why I suggested 2pi/360 instead of 180/pi like phi did.

OpenStudy (phi):

you just multiply 32 * 3.14159 then divide by 180 what do you get ?

OpenStudy (anonymous):

oh i see, thanks @whpalmer4

OpenStudy (whpalmer4):

When you actually calculate it, you can drop the common factor of 2 and use pi and 180

OpenStudy (anonymous):

@phi i got .56

OpenStudy (anonymous):

@whpalmer4 would that answer be correct for your way also?

OpenStudy (whpalmer4):

Yes, you should get the exact same answer.

OpenStudy (whpalmer4):

You are doing the same thing, after all :-)

OpenStudy (phi):

you should always get more decimals than what you round to. I get 0.5585 round to the nearest 100th place 0.56 so you got the right answer.

OpenStudy (anonymous):

thank you both so much, i dont know who to give best answer, i really appreciate it! could one of you please explain how i could find 170 degrees for the same type of problem? it would really help a lot

OpenStudy (phi):

here are some videos http://www.khanacademy.org/math/trigonometry/basic-trigonometry/radians_tutorial/v/introduction-to-radians There are also some for trig in general. If you eat popcorn while watching they are not bad.

OpenStudy (anonymous):

I just barely started trig, sorry I can't help. I'm only working with tan, cos, and sin, of a right triangle. but @phi seems to be helping enough(:

OpenStudy (phi):

do the same thing, except use 170 instead of 32

OpenStudy (whpalmer4):

\frac{x}{170} = \frac{2\pi}{360}\]\]\[360x=170*2\pi\]\[x=170*\pi/180\]

OpenStudy (whpalmer4):

Crud, the first line should have been \[\frac{x}{170}=\frac{2\pi}{360}\]

OpenStudy (whpalmer4):

It's just a simple proportion.

OpenStudy (anonymous):

if i cross multiply there what would i get

OpenStudy (anonymous):

i dont have a calculator on me it just ran out of batteries :o

OpenStudy (whpalmer4):

For back of the envelope work, 1 radian is about 57 degrees. Cross multiplying is what I did. \[360x = 170*2\pi\]

OpenStudy (phi):

you can use google. type 170*pi/180= into the google search window

OpenStudy (anonymous):

ooh okay i didn't know that thanks :) so 2.97?

OpenStudy (phi):

yes

OpenStudy (whpalmer4):

Sounds about right.

OpenStudy (anonymous):

thank you both so much, i will go over these both again to understand it better. you really helped me out a lot! :)

OpenStudy (phi):

also, take a peek at the videos

OpenStudy (anonymous):

what videos? :)